# VideoPoet: A Large Language Model for Zero-Shot Video Generation

### v3

@@ A.5.1 Additional Implementation Details
Since the first frame is tokenized separately, MAGVIT-v2 allows images to be represented in the same vocabulary as video. In addition to being more compact, images provide many learnable characteristics that are not typically represented in videos, such as strong visual styles (e.g., art paintings), objects which are infrequently seen in video, rich captions, and significantly more text-image paired training data. When training on images, we resize the images to 128$\times$128 which are then tokenized to a latent shape of $(1,16,16)$, or 256 tokens. We scale the MAGVIT-v2 model’s size and train it on the datasets discussed in Section 5.1. The training follows two steps: image training, inflation (Yu et al., 2023c) and video training. Due to images requiring fewer tokens, we can include roughly 5$\times$ more images per batch than videos, i.e. 256 image tokens vs. 1280 video tokens. We use up to a maximum of 64 text tokens for all of our experiments. For the <res> token, the resolution is only specified for $128\times 224$ output, $128\times 128$ resolution is assumed otherwise.
@@ A.5.2 Super-resolution implementation details
We use a 1B model for the first 2×2$\times$2 × spatial super-resolution stage and a 500M model for the second 2×2$\times$2 × stage. The first super-resolution stage models videos of $17\times 448\times 256$ pixels with a token sequence of shape $(5,56,32)$. The second stage models videos of $17\times 896\times 512$ pixels with a token sequence of shape $(5,112,64)$. The token sequences are obtained with the same MAGVIT-v2 (Yu et al., 2023c) tokenizer used for the base language model. The custom super-resolution transformer has local self-attention windows for vertical, horizontal and temporal layers of shape $(1,56,4),(1,8,32),(5,8,8)$ in the first stage and $(1,112,2),(1,4,64),(5,8,8)$ in the second stage, respectively (Fig. 3). The cross-attention layers attend to local windows in the low-resolution sequence isomorphic to self-attention windows but with half the spatial size.
We train the super-resolution stages on a dataset of 64M high-quality text-video pairs using the masked modeling objective of MAGVIT (Yu et al., 2023a), with token factorization into $k=2$ groups (Yu et al., 2023c). During inference, we use the sampling algorithm of MAGVIT-v2 (Yu et al., 2023c) with 24 sampling steps for each stage and classifier-free guidance scale (Ho & Salimans, 2022; Brooks et al., 2023) of $4.0/8.0$ for the text condition and $1.0/2.0$ for the low-resolution condition, in the first/second stage.
@@ A.5.3 Additional Evaluation details
A.5.4 Zero-shot Text-to-Video Evaluation Settings
@@
A.5.5 Self-Supervised Tasks Evaluation Settings
@@
A.5.6 Stylization Evaluation on DAVIS

### v4

@@ A.5.1 Additional Implementation Details
Since the first frame is tokenized separately, MAGVIT-v2 allows images to be represented in the same vocabulary as video. In addition to being more compact, images provide many learnable characteristics that are not typically represented in videos, such as strong visual styles (e.g., art paintings), objects which are infrequently seen in video, rich captions, and significantly more text-image paired training data. When training on images, we resize the images to 128$\times$128 which are then tokenized to a latent shape of $(1,16,16)$, or 256 tokens. We scale the MAGVIT-v2 model’s size and train it on the datasets discussed in Section 5.1. The training follows two steps: image training, inflation (Yu et al., 2024) and video training. Due to images requiring fewer tokens, we can include roughly 5$\times$ more images per batch than videos, i.e. 256 image tokens vs. 1280 video tokens. We use up to a maximum of 64 text tokens for all of our experiments. For the <res> token, the resolution is only specified for $128\times 224$ output, $128\times 128$ resolution is assumed otherwise.
@@ A.5.2 Super-resolution implementation details
We use a 1B model for the first 2×2$\times$2 × spatial super-resolution stage and a 500M model for the second 2×2$\times$2 × stage. The first super-resolution stage models videos of $17\times 448\times 256$ pixels with a token sequence of shape $(5,56,32)$. The second stage models videos of $17\times 896\times 512$ pixels with a token sequence of shape $(5,112,64)$. The token sequences are obtained with the same MAGVIT-v2 (Yu et al., 2024) tokenizer used for the base language model. The custom super-resolution transformer has local self-attention windows for vertical, horizontal and temporal layers of shape $(1,56,4),(1,8,32),(5,8,8)$ in the first stage and $(1,112,2),(1,4,64),(5,8,8)$ in the second stage, respectively (Fig. 3). The cross-attention layers attend to local windows in the low-resolution sequence isomorphic to self-attention windows but with half the spatial size.
We train the super-resolution stages on a dataset of 64M high-quality text-video pairs using the masked modeling objective of MAGVIT (Yu et al., 2023a), with token factorization into $k=2$ groups (Yu et al., 2024). During inference, we use the sampling algorithm of MAGVIT-v2 (Yu et al., 2024) with 24 sampling steps for each stage and classifier-free guidance scale (Ho & Salimans, 2022; Brooks et al., 2023) of $4.0/8.0$ for the text condition and $1.0/2.0$ for the low-resolution condition, in the first/second stage.
@@ A.5.3 Additional Evaluation details
A.5.4 Additional Human Evaluation details
Figure 8: Example screenshot of the user interface for human side-by-side comparisons.
Figure 8 shows an example screenshot of our side-by-side UI for comparing models. We used a team of 7 human raters to complete all ratings. To achieve best results, we call VideoPoet using the negative prompt “a still shot of an ugly cartoon, slideshow of an empty scene, low resolution, distorted and disfigured” and rewrite the given prompt by appending the string highly detailed, cinematic, arc shot, high contrast, soft lighting, 8k.
A.5.5 Zero-shot Text-to-Video Evaluation Settings
@@
A.5.6 Self-Supervised Tasks Evaluation Settings
@@
A.5.7 Stylization Evaluation on DAVIS